# Calibration of the strain amplitude recorded with DAS using a strainmeter array

### v1

@@ Abstract
The power of distributed acoustic sensing (DAS) lies in its ability to sample deformation signals along an optical fiber at hundreds of locations with only one interrogation unit (IU). While the IU is calibrated to record ‘fiber strain’, the properties of the cable and its coupling to the rock control the ‘strain transfer rate’ and hence how much of ‘rock strain’ is represented in the recorded signal. We use DAS recordings in an underground installation colocated with an array of strainmeters in order to calibrate the ‘strain transfer rate’ in situ, using earthquake signals between 0.05 Hz and 0.1 Hz. A tight-buffered cable and a standard loose-tube telecommunication cable (running in parallel) are used, where a section of both cables loaded down by loose sand and sand bags is compared to a section, where cables are just unreeled on the floor. The ‘strain transfer rate’ varies between 0.13 and 0.53 depending on cable and installation type. The sandbags show no obvious effect and the tight-buffered cable generally provides a larger ‘strain transfer rate’. Calibration of the ‘strain transfer rate’ with respect to the strainmeter does not depend on wave propagation parameters. Hence it is applicable to the large amplitude surface wave signal in a strain component almost perpendicular to the great-circle direction for which a waveform comparison with seismometer data does not work. The noise background for ‘rock strain’ in the investigated band is found at about an rms-amplitude of 0.1 nstrain in 1/6 decade for the tight-buffered cable. This allows a detection of marine microseisms at times of high microseism amplitude.
@@ 1 Introduction
Use cases, which rely on amplitude are less frequent. Such a use case would be the measurement of volume strain for the reduction of Newtonian Noise (Harms, 2015; Harms et al., 2022), as being planned for the Einstein Telescope (ET Steering Committee, 2020).
With well calibrated DAS strain recordings on cables running in six different directions from a single location, the full strain tensor could be composed. This requires a proper calibration of the ‘strain transfer rate’ (the fraction of ‘rock strain’ picked up as ‘fiber strain’) and a sufficiently low detection level, such that the strain background signal is resolved. Rademacher (2024) recently pointed at the limitations which still exist for DAS in both respects. We investigate these in the current study.
@@
Comparisons of DAS recordings with other types of observations are often carried out with respect to colocated seismometers. This approach involves translating strain or strain rate into displacement or velocity (or vice versa). This is only possible for non-dispersive plane waves with known ray-parameter. In this study, we directly compare the recordings of a DAS IU with those from the long-running array of Invar-wire strainmeters at the Black Forest Observatory (BFO). Both, the DAS recorded strain and the signal from the strainmeter can be expected to represent the same rock deformation, independent of its cause and independent of wave parameters in particular. Hence, this comparison allows for direct calibration of the DAS system’s amplitude response. Based on this calibration we estimate the detection threshold for strain recordings. The signal-to-noise ratio of both systems is appropriate for this purpose in the frequency band between 0.05 Hz and 0.1 Hz.
@@ 2.1 Distributed acoustic sensing (DAS)
Two fiber optic cables are installed in the tunnel. They are unreeled on the floor of the gallery and are in direct contact with the formation. In the section called ‘Anton Gang’ the cables are loaded down by sand and tightly spaced sandbags in order to improve the mechanical coupling to the rock, provide thermal shielding and cover the cable against water dripping from the gallery ceiling. The blue cable in Fig. 1 is a standard flexible telecommunication cable with loose-tube fibers embedded in gel. The green cable in Fig. 1 has a stiff jacket containing tight-buffered fibers (see section S1.2 in the supplemental material for additional details). We use two single mode fibers in each of the cables and splice them in series, such that the ‘Anton Gang’ and the ‘Vorstollen’ are both sampled four times. The signals from the fibers are sampled by a Febus A1-R interrogation unit (IU) installed in the laboratory (Fig. 1). The DAS recording parameters are listed in Table 1.
@@ 2.2 Reference instruments
BFO operates an array of three well calibrated, 10 m long, horizontal, Invar-wire strainmeters (Fig. 1). The SEED codes for the instruments are II.BFO.00.BSA, II.BFO.00.BSB, and II.BFO.00.BSC. Their design is based on the instruments by King and Bilham (1976) and are discussed in more detail by Zürn et al. (2015). Agnew (1986) and Zürn (2012) discuss instruments of this type and their properties. These instruments are primarily designed to record very-long period signals, such as tidal strain or Earth’s free oscillations. At short signal period they show a linear parasitic sensitivity to vertical ground acceleration because of the inertia of the pick-up system. They provide useful strain signals at frequencies below 1 Hz. We use the instruments at frequencies below 0.1 Hz (which provides a safety-margin to the parasitic response) as a reference to present ‘rock strain’.
@@ 3.1 Available data
From the GEOFON catalog (Quinteros et al., 2021) we find 84 earthquakes with moment magnitude larger than 6 in the recording time period. 21 of them show a maximum strain amplitude of larger than 1 nstrain in a visual inspection, which is considered large enough to provide a sufficient signal-to-noise ratio in the DAS data. For 19 of the events DAS data is available and allows an analysis. Event characteristics are detailed in the supplementary material (see Table S2). The largest amplitudes are found in the surface wave train. For body-waves, the horizontal strain amplitude is the smaller the steeper the ray incidence.
@@ 3.2 Pre-processing of DAS data
Signals sampled simultaneously on a section of the tight-buffered cable rolled up on a ‘Drum’ (see Fig. 1 and Table S1 in the supplemental material) get stacked and subtracted from the recordings in order to reduce common mode laser noise. For a gauge-length of 50 m this noise component clearly dominates the background level of the DAS signal at frequencies below 0.5 Hz and peaks at about 0.07 Hz with ‘fiber strain’-amplitudes of a few nstrain. By subtracting this coherent signal component, the background level is lowered by up to 20 dB near 0.1 Hz. Data is converted from strain rate to strain for which the noise floor rapidly increases at low frequencies, such that we focus our analysis on frequencies higher than 0.05 Hz.
@@ 4.1 Strain transfer rate
where is the DAS time series of ‘fiber strain’ and is the strainmeter time series of ‘rock strain’. Data are filtered by a Butterworth high-pass (0.05 Hz, 4th order) and low-pass (0.1 Hz, 4th order). The average is thus removed from the signals prior to the computation. Fig. 2 (left) shows the ‘strain transfer rate’ and Table 2 summarizes the ranges of values. They are between 0.13 and 0.53 in all cases and primarily depend on location and cable type. The scatter observed in the distribution for one cable at one location is smaller than the differences to the other combinations. The character of scatter of values does not significantly differ between cables or locations. Differences between colocated fibers in the same cable and location are smaller than 0.05 in most cases. We also find no discernible correlation of ‘strain transfer rate’ with signal amplitude, back-azimuth or other earthquake specific parameters (see Figs. S26 and S27 in the supplemental material for example). The largest values of ‘strain transfer rate’ are found for the tight-buffered cable, which was expected. Although the loose-tube cable in the ‘Anton Gang’ is loaded down by sandbags as well, its ‘strain transfer rate’ is the lowest among all fibers. Hence, we do not observe a significant benefit from sand and sandbags to improve overall coupling.
@@ 4.4 Comparison with seismometer data
where and are horizontal particle velocity and linear strain in azimuth of , respectively, is the azimuth of wave propagation, and is the horizontal component of slowness (see section S4.1 in the supplemental material for additional details). The accuracy of the scaled seismometer signal in representing strain depends on the validity of the plane-wave assumption underlying the conversion. While strainmeters and DAS can be expected to record the same quantity, namely linear strain in a given azimuth, independent of the nature of the signal, this is not the case for the seismometer derived strain signals.
@@
The surface waves with good signal-to-noise ratio used in the current study not only propagate at varying slowness (dispersion and superposition of Love- and Rayleigh-waves), they also contain significant non-plane wavefield components due to scattering in particular on continental paths. Wielandt (1993) discusses the consequences of the non-plane nature of waves for the spatial derivatives of displacement. The scaling in Eq. (7) lets vanish for being perpendicular to . In cases where comes close to being perpendicular to the actual strain component is dominated by off-great-circle propagation and non-plane wave contributions, which results in a large scatter of regression coefficients (see Fig. S32 in the supplemental material). In hese cases, seismometers cannot reasonably be used to calibrate DAS signals.
@@ 6 Conclusions
The waveform similarity of DAS recorded signals with strainmeter recorded signals as well as in an intercomparison of DAS signals is high for large amplitude signals. For the body- and surface waves recorded after the main shocks (Mw 7.7 and Mw 7.6) of the Kahramanmaraş earthquake sequence on February 6th 2023 the normalized correlation coefficient (NCC) with respect to the strainmeter is larger than 0.92 for all DAS fibers in the frequency band from 0.05 Hz to 0.1 Hz. In an intercomparison of colocated DAS fibers for the body-wave signals up to 1 Hz, the NCC is larger than 0.97 and reaches 1.000 in some cases. For smaller signal amplitude the DAS signal-to-noise ratio worsens and the NCC becomes smaller. The signal quality in this respect is better for the tight-buffered cable and worst for the loose-tube cable in the ‘Vorstollen’, where it is not protected by a sand bed against water dripping from the ceiling.
@@
The scatter of ‘strain transfer rate’ values for one cable at one location is smaller than the differences between cables and locations. We find no systematic dependency of ‘strain transfer rate’ on backazimuth of the event, nor do we find any other systematic correlation except with respect to cable type and location. The ‘strain transfer rate’ of the cables cannot be specified independently of local installation conditions.
@@ Acknowledgments
We are grateful to Gaetan Calbris and Vincent Lanticq from Febus Optics who provided helpful information on the measurement principle used in the Febus A1-R. Peter Duffner, Felix Bögelspacher, and Leon Merkel supported the installation of the experiment. The cable with the tight-buffered fibers (green cable) has been made available for the experiments through the INSIDE project. INSIDE is supported by the German Federal Ministry for Economic Affairs and Climate Action and the Project Management Jülich (PtJ) under the grant agreement number 03EE4008C.
@@
396.8 | loose-tube | 450102.6 | 5353139.4 |
620.0 | tight-buffered | 450098.3 | 5353146.5 |
1686.4 | tight-buffered | 450101.6 | 5353141.0 |
1909.6 | loose-tube | 450099.3 | 5353144.9 |
@@
297.6 | loose-tube | 450028.3 | 5353161.0 |
719.2 | tight-buffered | 450017.0 | 5353161.1 |
1587.2 | tight-buffered | 450023.4 | 5353161.1 |
2008.8 | loose-tube | 450022.0 | 5353161.1 |
@@
124.0 | tight-buffered | 450101.4 | 5353141.4 |
1190.4 | tight-buffered | 450098.5 | 5353146.2 |
1413.6 | loose-tube | 450102.4 | 5353139.7 |
@@
223.2 | tight-buffered | 450023.0 | 5353161.1 |
1091.2 | tight-buffered | 450017.4 | 5353161.1 |
1512.8 | loose-tube | 450027.9 | 5353161.0 |
2058.4 | loose-tube | 450019.3 | 5353161.1 |
@@
857.5 | tight-buffered, drum start | ||
1150.0 | tight-buffered, drum center | ||
1442.5 | tight-buffered, drum end | ||
@@
367.5 | tight-buffered, drum start | ||
660.0 | tight-buffered, drum center | ||
952.5 | tight-buffered, drum end |

### v2

@@ Abstract
The power of distributed acoustic sensing (DAS) lies in its ability to sample deformation signals along an optical fiber at hundreds of locations with only one interrogation unit (IU). While the IU is calibrated to record ‘fiber strain’, the properties of the cable and its coupling to the rock control the ‘strain transfer rate’ and hence how much of ‘rock strain’ is represented in the recorded signal. We use DAS recordings in an underground installation near an array of strainmeters in order to calibrate the ‘strain transfer rate’ in situ, using earthquake signals between 0.05 Hz and 0.1 Hz. A tight-buffered cable and a standard loose-tube telecommunication cable (running in parallel) are used, where a section of both cables loaded down by loose sand and sand bags is compared to a section, where cables are just unreeled on the floor. The ‘strain transfer rate’ varies between 0.13 and 0.53 depending on cable and installation type. The sandbags show no obvious effect and the tight-buffered cable generally provides a larger ‘strain transfer rate’. Calibration of the ‘strain transfer rate’ with respect to the strainmeter does not depend on wave propagation parameters. Hence it is applicable to the large amplitude surface wave signal in a strain component almost perpendicular to the great-circle direction for which a waveform comparison with seismometer data does not work. The noise background for ‘rock strain’ in the investigated band is found at about an rms-amplitude of 0.1 nstrain in 1/6 decade for the tight-buffered cable. This allows a detection of marine microseisms at times of high microseism amplitude.
@@ 1 Introduction
Use cases, which rely on amplitude are less frequent. Such a use case would be the measurement of volume strain for the reduction of Newtonian Noise (Harms, 2015; Harms et al., 2022), as being planned for the Einstein Telescope (ET Steering Committee, 2020). With well calibrated DAS strain recordings on cables running in six different directions from a single location, the full strain tensor could be composed. This requires a proper calibration of the ‘strain transfer rate’ (the fraction of ‘rock strain’ picked up as ‘fiber strain’) and a sufficiently low detection level, such that the strain background signal is resolved. Rademacher (2024) recently pointed at the limitations which still exist for DAS in both respects. We investigate these in the current study.
@@
The studies by Lindsey et al. (2020) and Paitz et al. (2020) are two examples for in-situ calibration experiments. Both use a Silixa iDAS IUs and both primarily use surface wave signals.
Lindsey et al. (2020) focus on the low-frequency band from about 0.08 Hz to 1 Hz. They not only describe the calibration experiment, but also provide a well-written introduction to DAS including the conversion from optical phase to strain, the coupling of rock to fiber, and typical optical sources of noise. The DAS signals in their study are recorded on a gel-filled loose-tube dark fiber installed in a conduit. They convert the signals to particle velocity based on the assumption of plane waves and compare with the recordings of a broad-band seismometer at 66 m distance. The needed scaling factor, which is phase velocity along the fiber, is obtained by fk-analysis of 1 km of DAS data, which allows to capture wave dispersion and does not rely on the assumption of great-circle propagation. Though the superposition of Rayleigh- and Love-waves (both propagating at difference phase velocity) might still present a problem. The authors use surface waves from four teleseismic earthquakes to demonstrate that the amplitude response of the DAS is like nominally expected for frequencies below 0.1 Hz, although there is a significant amplitude fluctuation of a factor of 10 along the fiber (Lindsey et al., 2020, their figure 6b). For frequencies higher than 0.1 Hz, they find DAS amplitudes larger than nominal by up to about 10 dB (about a factor of 3). The analysis of earthquakes and marine microseisms in this band yields inconsistent results in that the amplification seen for marine microseisms is larger. As a potential cause, the authors mention possible problems with the fk-analysis at near-perpendicular incidence of marine microseisms. While they prefer fiber coupling issues as the explanation for the amplification, they do not clarify how a passive mechanism could account for the signal amplification.
Paitz et al. (2020) derive a frequency dependent transfer function of DAS signals with respect to reference recordings for seven frequency bands, spanning a total range from 0.34 mHz to 60 Hz. As test signals, they use hydraulic stimulation and surface waves from four earthquakes (one per analyzed frequency band) and one icequake. The installation types of DAS cables in the various experiments differ significantly, without the possibility to track down systematic variations to the type of installation. The hydraulic stimulation experiment carried out for the lowest frequency bands is exceptional in that the DAS fiber is cemented in a borehole and in that a fiber-Bragg-grating (FBG) strainmeter could be used as a reference, while for the others strain had to be estimated from particle velocity. Though the results for this hydraulic stimulation does not appear exceptional in terms of deviation from the expected nominal response. Overall the authors report a variation of the amplitude response by 10 dB (about a factor of 0.3 to 3) around the nominal value.
When comparing DAS recordings with seismometer recordings a translating of strain or strain rate into particle displacement or velocity (or vice versa) is necessary. This is only possible for non-dispersive plane waves with known ray-parameter. The limitations of this approach and the resulting waveform dissimilarity are shown by Lindsey et al. (2020, their figure 4a) for example. Paitz et al. (2020) mention incorrect phase-velocity estimates as a possible source of inaccuracy.
In the current study, we directly compare the recordings of a DAS IU with those from the long-running array of Invar-wire strainmeters at the Black Forest Observatory (BFO). Both, the DAS recorded strain and the signal from the strainmeter can be expected to represent the same rock deformation, independent of its cause and independent of wave parameters in particular. Hence, this comparison allows for direct calibration of the DAS system’s amplitude response, not limited to a specific wave type and without additional uncertainties due to limited accuracy of ray-parameters. Based on this calibration we estimate the detection threshold for strain recordings. The signal-to-noise ratio of both systems is appropriate for this purpose in the frequency band between 0.05 Hz and 0.1 Hz.
Our study further extends beyond the work of Lindsey et al. (2020) and Paitz et al. (2020) by interrogating two fibers in each cable, allowing us to distinguish coupling issues that effect all fibers in a cable, from photonic effects (like optical fading), which are specific to each fiber. We use a significantly larger dataset, comprising 19 earthquakes, covering almost all backazimuths and include body waves in the analysis. The simultaneous interrogation of fibers in two cables at two locations enables us to investigate the direct differences observed under four different coupling conditions.
@@ 2.1 Distributed acoustic sensing (DAS)
Two fiber optic cables are installed in the tunnel. They are unreeled on the floor of the gallery and are in direct contact with the formation. In the section called ‘Anton Gang’ the cables are loaded down by sand and tightly spaced sandbags in order to improve the mechanical coupling to the rock, provide thermal shielding and cover the cable against water dripping from the gallery ceiling. The blue cable in Fig. 1 is a standard flexible telecommunication cable with loose-tube fibers embedded in gel. The green cable in Fig. 1 has a stiff jacket containing tight-buffered fibers (see section S1.2 in the supplemental material for additional details). We use two single mode fibers in each of the cables and splice them in series, such that the ‘Anton Gang’ and the ‘Vorstollen’ are both sampled four times. The signals from the fibers are sampled by a Febus A1-R interrogation unit (IU) installed in the laboratory (Fig. 1). The DAS recording parameters are listed in Table 1. The Febus A1-R recording time is synchronized by GPS, like all other digitizers involved in the current study. When the IU was installed in the mine, the GPS timing signal was provided through an optical link (Meinberg GOAL: GPS Optical Antenna Link).
@@ 2.2 Reference instruments
BFO operates an array of three well calibrated, 10 m long, horizontal, Invar-wire strainmeters (Fig. 1). The SEED codes for the instruments are II.BFO.00.BSA, II.BFO.00.BSB, and II.BFO.00.BSC. Their design is based on the instruments by King and Bilham (1976) and are discussed in more detail by Zürn et al. (2015). Agnew (1986) and Zürn (2012) discuss instruments of this type and their properties. These instruments are primarily designed to record very-long period signals, such as tidal strain or Earth’s free oscillations. For this reason they are equipped with an in-situ calibration device, which makes use of interferometrically calibrated ’Crapaudines’ (Verbaandert, 1959). The accuracy of this calibration for the strainmeters is about 2 per cent. Comparison against theoretical tidal strain ensures a stability of the calibration of about 5 per cent in the long run. Section S2.4 in the supplemental material gives further details.
At short signal period the Invar-wire strainmeters show a linear parasitic sensitivity to vertical ground acceleration because of the inertia of the pick-up system. They provide useful strain signals at frequencies below 1 Hz. We use the instruments at frequencies below 0.1 Hz (which provides a safety-margin to the parasitic response) as a reference to present ‘rock strain’.
@@ 3.1 Available data
From the GEOFON catalog (Quinteros et al., 2021) we find 84 earthquakes with moment magnitude larger than 6 in the recording time period. 21 of them show a maximum strain amplitude of larger than 1 nstrain in a visual inspection, which is considered large enough to provide a sufficient signal-to-noise ratio in the DAS data. For 19 of the events DAS data is available and allows an analysis. Their backazimuths (BAZ) cover all directions with an azimuthal gap of (BAZ between NE and NE). Event characteristics are detailed in the supplementary material (see Table S2). The largest amplitudes are found in the surface wave train. For body-waves, the horizontal strain amplitude is the smaller the steeper the ray incidence.
@@ 3.2 Pre-processing of DAS data
Waveforms sampled simultaneously on a section of the tight-buffered cable rolled up on a ‘Drum’ (see Fig. 1 and Table S1 in the supplemental material) get averaged and subtracted from the recordings in order to reduce common mode laser noise (Lindsey et al., 2020, their section 2.6 Optical Noise). For a gauge-length of 50 m this coherent noise component clearly dominates the background level of the DAS signal at frequencies below 0.5 Hz and peaks at about 0.07 Hz with ‘fiber strain’-amplitudes of a few nstrain. By subtracting this coherent signal component, the background level is lowered by up to 20 dB near 0.1 Hz. In the current application, it is essential that the drum is sufficiently decoupled from the rock and that the fiber coiled on the drum does not pick up the earthquake signal. Otherwise the correction procedure would affect the earthquake signal amplitude. In the supplemental material (section S2.3 and Figures S26 and S27) we demonstrate that no signature of the earthquake is apparent in the correction signal.
Data is converted from strain rate to strain for which the noise floor rapidly increases at low frequencies, such that we focus our analysis on frequencies higher than 0.05 Hz.
@@ 4.1 Strain transfer rate
where is the DAS time series of ‘fiber strain’ and is the strainmeter time series of ‘rock strain’. Data are filtered by a Butterworth high-pass (0.05 Hz, 4th order) and low-pass (0.1 Hz, 4th order). The average is thus removed from the signals prior to the computation. Fig. 2 (left) shows the ‘strain transfer rate’ and Table 2 summarizes the ranges of values. They are between 0.13 and 0.53 in all cases and primarily depend on location and cable type. The variability in the values calculated for a single installation (i.e., one cable at one location) is depicted using a violin plot, where the edges illustrate the density of dots in Fig. 2.
The variability within individual violin plots is smaller than the differences between the median values for different cables and locations. This observation remains consistent across all considered installations, regardless of cable type or location. It underlines the significance of the differences observed between the installations. Differences between colocated fibers for the same cable, location, and event are smaller than 0.05 in almost all cases, which lets us rule out that fiber related causes (like optical fading) would dominate the observed differences. We find no discernible correlation of ‘strain transfer rate’ with signal amplitude, back-azimuth or other earthquake specific parameters (see Figs. S28 and S29 in the supplemental material for example).
The largest values of ‘strain transfer rate’ are found for the tight-buffered cable, which was expected. Although the loose-tube cable in the ‘Anton Gang’ is loaded down by sandbags as well, its ‘strain transfer rate’ is the lowest among all fibers. Hence, we do not observe a significant benefit from sand and sandbags to improve overall coupling.
@@ 4.4 Comparison with seismometer data
where and are horizontal particle velocity and linear strain in azimuth of , respectively, is the azimuth of wave propagation, and is the horizontal component of slowness (see section S4.1 in the supplemental material for additional details). The accuracy of the scaled seismometer signal in representing strain depends on the validity of the plane-wave assumption underlying the conversion. While strainmeters and DAS can be expected to record the same quantity, namely linear strain in a given azimuth, independent of the nature of the signal, this is not the case for the seismometer derived strain signals. This is not only the case because the recorded waveform is a superposition of non-plane waves with different phase slowness, but also because of the local free surface affecting the particle velocity and the strain in different ways (see section S2.5 in the supplemental material for additional details).
@@
The surface waves with good signal-to-noise ratio used in the current study not only propagate at varying slowness (dispersion and superposition of Love- and Rayleigh-waves), they also contain significant non-plane wavefield components due to scattering in particular on continental paths. Wielandt (1993) discusses the consequences of the non-plane nature of waves for the spatial derivatives of displacement. The scaling in Eq. (7) lets vanish for being perpendicular to . In cases where comes close to being perpendicular to the actual strain component is dominated by off-great-circle propagation and non-plane wave contributions, which results in a large scatter of regression coefficients (see Fig. S34 in the supplemental material). In these cases, seismometers cannot reasonably be used to calibrate DAS signals.
@@ 6 Conclusions
The waveform similarity between DAS- and strainmeter-recorded signals, as well as in the intercomparison of DAS signals, is high for signals with large amplitudes. For the body- and surface waves recorded after the main shocks (Mw 7.7 and Mw 7.6) of the Kahramanmaraş earthquake sequence on February 6th 2023 the normalized correlation coefficient (NCC) with respect to the strainmeter is larger than 0.92 for all DAS fibers in the frequency band from 0.05 Hz to 0.1 Hz. In an intercomparison of colocated DAS fibers for the body-wave signals up to 1 Hz, the NCC is larger than 0.97 and reaches 1.000 in some cases. For smaller signal amplitude the DAS signal-to-noise ratio worsens and the NCC becomes smaller. The signal quality in this respect is better for the tight-buffered cable and worst for the loose-tube cable in the ‘Vorstollen’, where it is not protected by a sand bed against water dripping from the ceiling.
@@
The scatter in ‘strain transfer rate’ values for a single cable at one location is smaller than the differences observed between different cables and locations. By the high similarity of signals from two fibers in the same cable, we can rule out fiber related causes for the difference, like optical fading. We find no systematic dependency of ‘strain transfer rate’ on backazimuth of the event, nor do we find any other systematic correlation except with respect to cable type and location. The ‘strain transfer rate’ of the cables cannot be specified independently of local installation conditions.
@@ Acknowledgments
We thank Peter Duffner for taking care of the BFO strainmeters and their calibration devices. We are grateful to Walter Zürn for many fruitful discussions and his continuing support of the operation of the Invar-wire strainmeters. Gaetan Calbris and Vincent Lanticq from Febus Optics provided helpful information on the measurement principle used in the Febus A1-R. Peter Duffner, Felix Bögelspacher, and Leon Merkel supported the installation of the experiment. We are grateful for the well-considered comments provided by two anonymous reviewers. The cable with the tight-buffered fibers (green cable) has been made available for the experiments through the INSIDE project. INSIDE is supported by the German Federal Ministry for Economic Affairs and Climate Action and the Project Management Jülich (PtJ) under the grant agreement number 03EE4008C.
@@
| 396.8 | loose-tube | 450102.6 | 5353139.4 |
| 620.0 | tight-buffered | 450098.3 | 5353146.5 |
| 1686.4 | tight-buffered | 450101.6 | 5353141.0 |
| 1909.6 | loose-tube | 450099.3 | 5353144.9 |
@@
| 297.6 | loose-tube | 450028.3 | 5353161.0 |
| 719.2 | tight-buffered | 450017.0 | 5353161.1 |
| 1587.2 | tight-buffered | 450023.4 | 5353161.1 |
| 2008.8 | loose-tube | 450022.0 | 5353161.1 |
@@
| 124.0 | tight-buffered | 450101.4 | 5353141.4 |
| 1190.4 | tight-buffered | 450098.5 | 5353146.2 |
| 1413.6 | loose-tube | 450102.4 | 5353139.7 |
@@
| 223.2 | tight-buffered | 450023.0 | 5353161.1 |
| 1091.2 | tight-buffered | 450017.4 | 5353161.1 |
| 1512.8 | loose-tube | 450027.9 | 5353161.0 |
| 2058.4 | loose-tube | 450019.3 | 5353161.1 |
@@
| 857.5 | tight-buffered, drum start | ||
| 1150.0 | tight-buffered, drum center | ||
| 1442.5 | tight-buffered, drum end | ||
@@
| 367.5 | tight-buffered, drum start | ||
| 660.0 | tight-buffered, drum center | ||
| 952.5 | tight-buffered, drum end | ||
@@
S2.3 Reduction of coherent noise
The Febus A1-R recordings contain a significant component of coherent noise at frequencies below 0.5 Hz. Lindsey et al. (2020, their section 2.6 Optical Noise) attribute this type of noise to vibrations of the optoelectronic system.
Other studies use the average over all recorded channels to capture the coherent component of noise. This relies on the earthquake signal not being coherent (because of propagation delay and varying orientation of the fiber), such that it more or less cancels out in the average and the earthquake signal in a single channel is not deteriorated by subtracting the average over all channels.
We follow a different approach. The idea behind using the reference coil on the drum is, that this section of the fiber is sufficiently decoupled from the rock and hence does not pick up a signal of the teleseismic earthquakes. For this reason it can be subtracted without affecting the amplitude of the earthquake signal.
In Figs. S26 and S27 we demonstrate this with the recordings of the Mw 7.7 Pazarcık earthquake. This is the earthquake which produced the by far largest strain amplitudes in the set of analyzed earthquakes. Even for the large amplitude surface waves, no signature of the earthquake signal is present in the recording from the reference coil (center diagram in Fig. S27). If a fraction of the earthquake signal should be hidden in the coherent noise on the reference drum, its amplitude is reduced to below 1.5 per cent of the full signal.
The P-wave signal of the earthquake becomes only apparent after this type of correction is applied, as shown in Fig. S26.
S2.4 Calibration of the strainmeters
The strainmeters operated at BFO are Invar-wire strainmeters of the design by King and Bilham (1976). Agnew (1986) and Zürn (2012) discuss instruments of this type and their properties. These instruments are primarily designed to record very-long period signals, such as tidal strain or Earth’s free oscillations. For the analysis of tidal strain amplitudes instrument calibration is of utmost importance. The original calibration of the instruments was based on separate calibrations of several components, which then were put in series. In unfortunate cases the individual calibration inaccuracies added up constructively, which resulted in an unacceptable inaccuracy for the entire strainmeter.
For this reason an in-situ calibration mechanism for these instruments was developed at BFO, based on so-called ’Crapaudines’, which provides an accuracy of about 2 per cent. They were originally used by Verbaandert (1959, 1962, 1963) in the calibration of tiltmeters. This application is described by Melchior (1966, pages 161 to 166), Melchior (1978, his section 8.8) and Agnew (1986, his section 4.1.5). Melchior (1966) calls them ’Expandable bearing plates of Verbaandert’. Agnew (1986) calls them ’distensible support of Verbaandert’. The ’Crapaudines’ themselves are calibrated interferometrically (Verbaandert, 1959, 1962) and are used to impose a well defined motion on the fixed end of the wire in the strainmeter. The calibration factors of a few ’Crapaudines’ used at BFO were confirmed by an independent interferometric measurement.
Because calibrations disturb the long-term tidal recordings, the stability of the calibration is occasionally tested by a comparison against theoretical tidal strain (Longman, 1959). This way we ensure a calibration accuracy and stability of about 5 per cent. Zürn et al. (2015) documented an episode of unstable calibration during the 2011 Tōhoku earthquake and how this was resolved. They were able to obtain strain amplitudes for radial mode 0S0, which were consistent with Earth’s radius and the observed radial surface motion within 15 per cent after the 2004 Sumatra-Andaman earthquake (their table 2).
The STR cannot be more accurate than the strainmeter calibration. The variation of STR between locations and cable type, however, cannot be due to limited calibration accuracy. This is because the differences exceed the level of 5 per cent and because the STR for the blue cable is larger in the ‘Anton Gang’ than in ‘Vorstollen’, while it is the opposite for the green cable, which cannot be caused by a calibration bias in principle.
S2.5 Distortion of the strain field due to local heterogeneity
Strain measurements and tilt measurements both are affected by local heterogeneity, sometimes called ’cavity effects’. This was first reported by King and Bilham (1973) and Baker and Lennon (1973) for measurements of tidal tilt and limits the usefulness of tidal strain and tilt amplitudes in the investigation of Earth’s elastic properties. These effects are strongest at free surfaces of the chambers (cavities) in which instruments are installed and near the free surface of the Earth, which can have a significant topography. The free-surface condition lets components of the stress tensor vanish and this enforces specific linear combinations for the components of the deformation tensor through the stress-strain relation. Harrison (1976) investigated this by finite-element modeling.
All strain measurements are affected by such strain-strain coupling. This applies to DAS recordings as well as to conventional strainmeters. In a more general sense the coupling issues between rock and fiber could also be summarized under these effects of local heterogeneity. Likewise particle motion is affected by free-surface conditions (Kennett, 1991). When comparing DAS recorded strain with signals from reference instruments, these effects generally could limit the usefulness of the observations, but are commonly ignored in the respective studies (except for cable coupling). The problem is even more critical, when comparing DAS recorded strain with particle motion recorded by seismometers, because the effect of the free surface on both is different and most likely the plane wave assumption (see section S4 ‘Plane wave strain from seismometer recordings’) is violated in this way. If DAS recorded strain is compared to recordings of a strainmeter, as is done in the current study, a problem may arise from spatially varying strain-strain coupling, if instruments are installed at a distance from each other.
The 10 m long Invar-wire strainmeters show little cavity effects due to their installation in the center and along the symmetry axes of 60 m long tunnels. For the DAS cables the cavity effects in a more general sense are addressed as coupling issues controlled by installation and cable type.
Topography at BFO, however, alters the strain field as demonstrated by Emter and Zürn (1985) with finite element simulations. In particular the strainmeter in azimuth NE experiences significant strain-strain coupling through the local topography. Zürn et al. (2015, their table S2 in the supporting material) list the strain-strain coupling coefficients derived from tidal analysis. Their analysis of the radial mode 0S0 (Zürn et al., 2015, their table 2) suggests that not the same coupling might be at work for this type of straining. The factor 0.58 (for NE) given by Zürn et al. (2015, their table S2 in the supporting material) for tidal analysis is consistent with the factor of 0.67 derived in a 2D finite element analysis by Emter and Zürn (1985, their figure 5). For strain in the azimuth NE of the ‘Vorstollen’ we presumably observe part of this topography effect. The seismometer data systematically over-estimates the strain amplitude in this azimuth, as discussed in section S4.2 ‘Regression with respect to seismometer data’.
In the current experiment we use the Invar-wire strainmeter in a comparison with DAS data recorded at a distance of about 350 m. If the strain-strain coupling would vary along this distance, it would contribute to the amplitude difference seen between DAS recordings and strainmeter recordings. The difference shown by Emter and Zürn (1985, their Fig. 5), with strain amplitude increasing towards the tunnel entrance near the free surface, however, is opposite to the observed difference. Because the section of the ‘Vorstollen’, which is used in the current analysis, is at distance of about 200 m to the entrance the expected strain-strain coupling would be almost the same as for the strainmeters, according to Emter and Zürn (1985, their Fig. 5).
For strain recordings with cemented fibers in the ‘Vorstollen’ Forbriger et al. (2024) recently found a STR of about 1 with respect to the strainmeters for four different IUs, independently. We take this as a confirmation that the DAS cable and the strainmeters experience the same topography-effect, indeed.
@@
Only after subtracting the signal from the reference coil, the P-wave arrival (at 70 s on the time scale) becomes apparent.
The surface waves of the Mw 7.7 Pazarcık earthquake are the largest amplitude strain signals in the entire analyzed data. No signature of this earthquake signal, which has a signal-to-noise ratio of about 60 with respect to the coherent noise in the raw recording shown in the top panel, is apparent in the recording from the reference drum.